A source of sound vibrates according to the equation y = 0.05 Cos π t. It sends out eaves of velocity 1.5ms-1. The wavelength of the waves is  (a) 1.5 m     (b) 3.0     (c) 4.5 m      (d) 6.0 m   
 

Grade:12th pass

1 Answers

Aditya Narayan
5 years ago
wave velocity is, 
w/k=1.5
the general equation is Asin(wt ± kx) 
w=π
k=2π/λ
π/2π/λ=1.5
λ/2=1.5
so, λ=3